OpenStudy (anonymous):

without using your calculator, find the values of the 6 trigonometric functions for each angle.

OpenStudy (anonymous):

1: 315 degrees 2: 270 degrees 3: \[-\frac{ 2\pi }{ 3 }radians\] 4:\[\frac{ 5\pi }{ 6 }radians\]

OpenStudy (anonymous):

@vera_ewing lol

OpenStudy (anonymous):

:P help?

OpenStudy (lynfran):

use the unit circle

OpenStudy (anonymous):

ok but how?

OpenStudy (anonymous):